# ContextCache: Context-Aware Semantic Cache for Multi-Turn Queries in Large Language Models

### v2

@@ 2. BACKGROUND
Semantic caching extends beyond traditional keyword-based caching by operating on meaning rather than lexical similarity. This approach enables systems to identify semantic equivalence between linguistically diverse queries, facilitating response reuse when similar questions appear in different lexical forms.
Embedding generation. The embedding generator forms the core component of semantic caching systems, transforming queries into dense vector representations in high-dimensional space. Modern embedding systems employ transformer-based (Vaswani et al., 2017) neural architectures with multi-head self-attention mechanisms to capture contextual relationships between tokens. These models (Lan et al., 2019) encode semantic information into vectors where geometric proximity corresponds to semantic similarity between the original text inputs. This transformation can be represented as: $E(q)=f_{\theta}(q)$, where $E(q)$ is the embedding of query $q$, and $f_{\theta}$ denotes the transformer-based embedding function with parameters $\theta$.
Similarity quantification. Semantic similarity evaluation provides the decision mechanism for the caching framework, typically implemented through cosine similarity between vector representations. This metric measures the angular proximity between embeddings, formulated as: $\text{similarity}(E_{1},E_{2})=\frac{E_{1}\cdot E_{2}}{||E_{1}||\,||E_{2}||}$, where $E_{1}$ and $E_{2}$ represent embedding vectors, $E_{1}\cdot E_{2}$ denotes their dot product, and $||E_{1}||$ and $||E_{2}||$ represent their respective Euclidean norms. When similarity exceeds a predetermined threshold, the system classifies queries as semantically equivalent, retrieving cached responses instead of invoking the LLM.
@@ 4. DEMONSTRATION OVERVIEW
Demonstration engagement. To quantify system effectiveness, we integrated an interactive dashboard (Figure 5) that visualizes ContextCache’s operational metrics in real-time. The interface tracks key performance indicators, including Cache Positive/Negative Hits, LLM API Calls, Query Volume, and Memory Utilization. The execution log provides request-level performance data, enabling users to examine query processing paths and response timing. Through this log analysis, users can clearly observe that cache-served responses deliver approximately 10 times lower latency compared to LLM invocation, while maintaining high semantic matching precision in multi-turn dialogues. This demonstrates that ContextCache is an efficient solution for reducing computational costs and response latency in LLM applications.

### v3

@@ 2. BACKGROUND
Semantic caching improves upon traditional methods by operating on a query’s meaning rather than its exact words. This allows the system to identify semantically equivalent questions, even when phrased differently, and reuse answers for greater efficiency.
Embedding generation. The core of the system is the embedding generator, which transforms queries into vector representations. It employs transformer-based architectures (Vaswani et al., 2017) to encode semantic information: $E(q)=f_{\theta}(q)$, where geometric proximity in the vector space corresponds to semantic similarity (Lan et al., 2019).
Similarity quantification. This step provides the decision mechanism for the cache. It uses cosine similarity to measure the angular proximity between two embedding vectors, formulated as: $\text{similarity}(E_{1},E_{2})=\frac{E_{1}\cdot E_{2}}{||E_{1}||\,||E_{2}||}$. When this similarity value exceeds a predetermined threshold, the system classifies queries as equivalent and retrieves cached response instead of invoking the LLM.
@@ 4. DEMONSTRATION OVERVIEW
Demonstration engagement. To quantify system effectiveness, we integrated an interactive dashboard (Figure 5) that visualizes ContextCache’s operational metrics in real-time. The interface tracks key performance indicators, including Cache True/False Hits, LLM API Calls, Query Volume, and Memory Utilization. The execution log provides request-level performance data, enabling users to examine query processing paths and response timing. Through this log analysis, users can clearly observe that cache-served responses deliver approximately 10 times lower latency compared to LLM invocation, while maintaining high semantic matching precision in multi-turn dialogues. This demonstrates that ContextCache is an efficient solution for reducing computational costs and response latency in LLM applications.